### v1

@@ 2.6 PathSegmentor for Explainable Cancer Diagnosis
Explainable AI in medical image analysis is essential for enhancing the interpretability and trustworthiness of diagnostic models [34, 35, 36]. In pathology, accurate cancer diagnosis relies on the identification, morphological assessment, and quantitative analysis of specific pathological objects, such as characteristic tissue patterns, cell structures, and nuclear features [37, 38]. Therefore, precise segmentation of these objects is critical for diagnostic decision-making. By extracting and analyzing diagnostically relevant biomarkers, PathSegmentor can generate interpretable, human-understandable explanations that help pathologists make more precise, evidence-based diagnostic judgments.
@@ 2.6.1 Classification-Segmentation for Feature Importance Estimation
The standard classification model for breast cancer diagnosis is trained on WSIs from the TCGA-BRCA dataset, following the conventional pipeline of patch feature extraction, slide aggregation, and final classification (Fig. 11a and Methods). This model achieves high classification performance with a macro AUC of 0.936 (Supplementary Table 19). To investigate model behavior, we implement an object-based perturbation approach to estimate feature importance. Different from traditional methods that use non-semantic perturbations, such as gray squares [41] or superpixel occlusions [42], our approach aims to perform biologically meaningful perturbations. We utilize PathSegmentor to segment predefined pathological objects within an image, and then apply localized blurring perturbations exclusively to these segmented regions. Feature importance is quantified by measuring the relative increase in prediction error (Methods).
For example, when applying perturbations to segmented regions of breast-tissue-tumor, the model’s prediction error increases from 0.003 to 0.012 (Fig. 10a, top right). Given a set of predefined pathological objects relevant to IDC and ILC diagnosis, the analysis across all TCGA-BRCA testing samples (Fig. 10a, bottom) reveals that these pathological objects generally contribute more significantly to the IDC class than ILC. Notably, the highly discriminative IDC features discovered by PathSegmentor, including breast-nuclei-ductal epithelium, breast-cell-cancerous, breast-nuclei-neoplastic, breast-tissue-tumor, breast-nuclei-epithelial, and breast-tissue-dcis, well correspond to established pathological diagnostic criteria for the IDC class [43]. These results establish PathSegmentor as an effective explanation tool that both validates model’s learning of medically relevant patterns and provides interpretable insights by quantifying decisive features in histological terms.
@@ 2.6.2 Segmentation-Classification for Imaging Biomarker Discovery
Another approach to improve the interpretability of standard classification models for cancer diagnosis is to develop an object-aware classification model capable of generating object-aware visual explanations. In our experiments, we build a segmentation-classification framework (Fig. 10b), first leveraging PathSegmentor to predict masks of pathological objects, and then decomposing whole-image features into explicit pathological object representations, using object-based multi-instance learning to make final predictions (Fig. 11b and Methods). While Class Activation Maps (CAMs) [44] from standard classification models highlight prediction-relevant regions that potentially serve as imaging biomarkers [45], their clinical utility is limited by an inability to specify the pathological nature of salient regions [46]. Our object-aware classification model bridges this semantic gap by generating object-aware CAMs, which not only localize discriminative regions but also associate them with specific pathological objects. This capability significantly enhances clinical interpretability, as pathologists require both precise pathological identification and spatial localization to discover biomarkers and inform diagnostic decisions.
After training on the TCGA-BRCA dataset, our object-aware classification model maintains diagnostic performance with a macro AUC of 0.953 for IDC and ILC classification (Supplementary Table 19). We generate object-aware CAMs by measuring the similarity between object features and classifier weights (Methods). Comparing vanilla CAMs with object-aware CAMs (Fig. 10c), we observe that both methods identify similar discriminative regions indicated by warmer color intensities, while our method additionally provides semantic information about these regions. For the IDC case in the top row, our object-aware CAM accurately localizes compact tumor mass regions and, importantly, identifies imaging biomarkers with high activation values that include breast-tissue-necrosis or debris, breast-tissue-tumor, breast-nuclei-epithelial. For the ILC case in the bottom row, our method effectively captures dispersed infiltration patterns of cancerous cells and ranks the salience of pathological biomarkers, highlighting the top three as breast-cell-cancerous, breast-nuclei-ductal epithelium, breast-nuclei-neoplastic. By resolving the semantic ambiguity of traditional CAMs, PathSegmentor delivers clinically meaningful visual explanations that assist pathologists in their diagnostic processes and has the potential to uncover previously under-recognized biomarker candidates.
@@ 3 Discussion
2) PathSegmentor could benefit from integrating universal prompts. PathSegmentor’s text prompts enable efficient segmentation; however, performance may decline with semantically novel categories. Spatial prompts provide complementary localization precision. We will combine multiple prompts [14, 48] to enhance robustness, especially for unseen categories in complex scenarios.
3) PathSegmentor has not yet undergone extensive clinical validation. The clinical utility of PathSegmentor requires validation through multicenter trials that assess real-world segmentation accuracy and robustness. Iterative refinements of PathSegmentor will be guided by pathologist feedback to ensure seamless integration into clinical workflows.
@@ 4.1 PathSeg Dataset Construction
In this work, we create the PathSeg dataset by integrating 17 publicly available pathology image segmentation datasets, totaling 275k image-mask-label triples. The statistical details and download links of each dataset are presented in Supplementary Table 1. The diverse pathology image samples included in the PathSeg dataset cover 20 anatomical regions, 3 histological structures, and 61 object types.
However, integrating these datasets for training presents two key challenges. 1) Ambiguous labels. Original labels typically specify only object types while neglecting critical anatomical and histological context. For instance, identically labeled cells from different anatomical regions exhibit distinct morphological characteristics, yet current labels fail to capture these biologically meaningful differences. Besides, identical labels may describe distinct biological entities across datasets, such as tumor variably denotes either tumor tissue or tumor cells. This definitional ambiguity critically necessitates standardized label hierarchies that explicitly distinguish between anatomical regions and histological structures. 2) Heterogeneous data. Pathology image datasets exhibit substantial variations in image characteristics, including scales of patch or whole-slide images, image resolutions ranging from $256\times 256$ to $5412\times 7215$, and magnification levels of 100$\times$, 40$\times$, 20$\times$, 10$\times$. Such inconsistencies prevent the model from learning unified feature representations. In the following parts, we present details of label reformat and data standardization.
@@
Data standardization. Prior to model training, we perform data preprocessing to standardize images from different datasets. The processing includes the following steps on both images and corresponding masks. 1) Magnification normalization. According to dataset metadata, images from 11 datasets in the PathSeg dataset have an initial magnification of 40×40$\times$40 × (Supplementary Table 1). To standardize magnification, images from the remaining datasets are rescaled to 40×40$\times$40 × using bilinear interpolation, while the corresponding masks are rescaled using nearest-neighbor interpolation. This approach ensures that the same object type maintains a relatively consistent size across different images, preventing significant variations in appearance. 2) Patching. Whole-slide images are split into a series of image patches using a sliding window. For any image dimension $D$ in height or width, if it exceeds 1500 pixels, we apply a sliding window of size 1024. The uniform overlap between adjacent patches is calculated as $\frac{(1024\times\left\lceil D/1024\right\rceil)-L}{\left\lceil D/1024\right\rceil-1}$. This ensures complete coverage while maintaining spatial regularity. The same method is applied to corresponding masks. 3) Resolution standardization. Image patches are resized to a consistent 1024$\times$1024 resolution via bilinear interpolation, and the mask patches are resized using nearest-neighbor interpolation. This unified input dimension facilitates batch processing during training. Finally, we split the datasets for model training and testing (Supplementary Table 1). We use the official split when available. If no official split is provided, we randomly divide the original dataset into 80% for training and 20% for testing.
@@ 4.2 PathSegmentor Architecture
[…]
@@
[…]
@@
Second, the image-enhanced queries $\mathbf{q}^{\prime}\in\mathbb{R}^{n\times d}$ are concatenated with text features $F_{\text{text}}$ to perform multi-head self-attention. Specifically, $[\mathbf{q}^{\prime}\,||\,F_{\text{text}}]\in\mathbb{R}^{(n+L)\times d}$ serve as the query, key, and value for computing the joint features $F_{\text{joint}}$, formulated as:
[…]
@@
Objective function. We optimize the model using a loss function combining binary cross-entropy (BCE) and Dice loss. For a ground-truth mask $\textbf{y}\in\{0,1\}^{H\times W}$ and predicted mask $\hat{\textbf{y}}\in\{0,1\}^{H\times W}$, the per-sample loss is computed as:
@@ 4.4 Implementation Details
Explainable cancer diagnosis. For both standard and object-based cancer diagnosis models (Fig. 11), we employ the image encoder of CONCH [26] as a frozen feature extractor, training only the slide aggregator and classifier. We adopt Attention-Based Multiple Instance Learning (ABMIL) [54] as the slide aggregator and a fully connected layer as the classifier. Following the CLAM [55] toolkit’s preprocessing pipeline, we extract patches with 1024$\times$1024 pixels at level 0 for all slides, resizing them to 224$\times$224 pixels while excluding slides with insufficient tissue area. We use the Adam [56] optimizer with a learning rate of 2e-4 and a weight decay of 1e-5. The diagnosis models are implemented in PyTorch [53] and run on one NVIDIA GTX3090 GPU.
@@ 4.6 Competing Methods
All datasets in the PathSeg dataset are publicly available, with a full list of dataset links and statistical details being presented in Extended Data Table 1.
@@
The code for PathSegmentor will be released at https://github.com/zhi-xuan-chen/PathSegmentor. For the competing methods, we adopt the official implementations of nnU-Net (https://github.com/MIC-DKFZ/nnUNet), DeepLabV3+ (https://github.com/VainF/DeepLabV3Plus-Pytorch), SAM-Path (https://github.com/cvlab-stonybrook/SAMPath), MedSAM (https://github.com/bowang-lab/MedSAM), SAM-Med2D (https://github.com/OpenGVLab/SAM-Med2D), and BiomedParse (https://github.com/microsoft/BiomedParse).
@@ Acknowledgment
Appendix A Extended Data
Table 1: An overview of the 17 data sources in our PathSeg dataset for pathology image segmentation. The table lists dataset name, magnification (MA), anatomical regions (AR), number of labels, images, masks (training/testing masks), and download links.
@@ Appendix A Extended Data
BCSS1  [60] 40x Breast 19 2,914 9,192 (7,322/1,870)
CoCaHis2  [61] 40x Liver 2 82 82 (58/24)
CoNIC20223  [62] 20x Colon 7 4,841 19,332 (15,387/3,945)
CRAG4  [63] 20x Colon 2 1,750 1,750 (1,429/321)
DigestPath20195  [64] 40x Colon 2 13,391 13,391 (10,725/2,666)
GlaS6  [65] 20x Colon 2 165 165 (85/80)
Janowczyk7  [66] 40x Breast 2 500 500 (400/100)
MIDOG8  [67] 40x Unspecified 3 1,482 1,802 (1,448/354)
MoNuSAC9  [68] 40x Unspecified 5 359 642 (454/188)
NuCLS10   [69] 40x Breast 13 1,612 4,443 (3,538/905)
PanNuke11   [70] 40x Unspecified 86 7,558 16,223 (13,017/3,206)
RINGS12   [71] 100x Prostate 3 38,227 43,739 (29,451/14,288)
SegPath13   [72] 40x Unspecified 9 133,645 133,645 (122,536/11,109)
SICAPv214   [73] 20x Prostate 2 23,924 23,924 (19,140/4,784)
TNBC15   [74] 40x Breast 12 66 167 (134/33)
WSSS4LUAD16   [75] 10x Lung 4 2,863 5,735 (4,590/1,145)
CoNSeP17   [8] 40x Colon 7 40 121 (0/121)
@@ Appendix B More Results in Internal Evaluation
GlaS 0.735 [0.702, 0.765] 0.833 [0.816, 0.850] 0.905 [0.885, 0.922] 0.843 [0.814, 0.868] 0.650 [0.621, 0.675] 0.862 [0.842, 0.880] 0.937 [0.924, 0.947] $\leq 0.01$
CRAG 0.752 [0.733, 0.772] 0.814 [0.798, 0.828] 0.895 [0.878, 0.911] 0.765 [0.746, 0.783] 0.583 [0.561, 0.604] 0.559 [0.531, 0.588] 0.933 [0.925, 0.940] $\leq 0.01$
WSSS4LUAD 0.677 [0.661, 0.692] 0.755 [0.742, 0.767] 0.838 [0.825, 0.850] 0.405 [0.391, 0.419] 0.379 [0.365, 0.392] 0.036 [0.032, 0.039] 0.833 [0.822, 0.843] -
CoCaHis 0.663 [0.582, 0.736] 0.757 [0.704, 0.806] 0.822 [0.755, 0.881] 0.611 [0.518, 0.692] 0.509 [0.388, 0.609] 0.679 [0.615, 0.734] 0.827 [0.791, 0.861] $0.06$
SICAPv2 0.608 [0.597, 0.618] 0.682 [0.674, 0.691] 0.622 [0.612, 0.632] 0.806 [0.801, 0.811] 0.700 [0.694, 0.706] 0.241 [0.235, 0.246] 0.821 [0.813, 0.828] $\leq 0.01$
RINGS 0.602 [0.596, 0.608] 0.733 [0.728, 0.739] 0.661 [0.655, 0.667] 0.775 [0.771, 0.779] 0.585 [0.582, 0.589] 0.246 [0.243, 0.248] 0.800 [0.794, 0.805] $\leq 0.01$
DigestPath2019 0.643 [0.631, 0.656] 0.539 [0.525, 0.554] 0.661 [0.650, 0.674] 0.764 [0.757, 0.772] 0.648 [0.640, 0.657] 0.508 [0.498, 0.518] 0.784 [0.775, 0.794] $\leq 0.01$
CoNIC2022 0.461 [0.452, 0.470] 0.360 [0.351, 0.368] 0.650 [0.642, 0.658] 0.202 [0.194, 0.211] 0.169 [0.163, 0.176] 0.287 [0.279, 0.296] 0.681 [0.674, 0.689] $\leq 0.01$
MoNuSAC 0.534 [0.493, 0.573] 0.493 [0.456, 0.533] 0.644 [0.605, 0.682] 0.359 [0.323, 0.400] 0.347 [0.310, 0.387] 0.288 [0.240, 0.336] 0.657 [0.615, 0.698] $\leq 0.01$
PanNuke 0.541 [0.530, 0.551] 0.336 [0.324, 0.347] 0.564 [0.553, 0.574] 0.364 [0.355, 0.374] 0.307 [0.298, 0.317] 0.621 [0.612, 0.631] 0.631 [0.621, 0.641] $\leq 0.01$
BCSS 0.393 [0.377, 0.408] 0.461 [0.445, 0.478] 0.506 [0.488, 0.523] 0.568 [0.555, 0.581] 0.508 [0.496, 0.521] 0.126 [0.117, 0.135] 0.620 [0.605, 0.636] $\leq 0.01$
SegPath 0.162 [0.157, 0.167] 0.104 [0.100, 0.108] 0.122 [0.117, 0.126] 0.215 [0.210, 0.219] 0.194 [0.190, 0.199] 0.083 [0.080, 0.085] 0.442 [0.436, 0.447] $\leq 0.01$
TNBC 0.209 [0.130, 0.296] 0.251 [0.137, 0.379] 0.473 [0.350, 0.593] 0.260 [0.184, 0.348] 0.229 [0.160, 0.306] 0.361 [0.243, 0.480] 0.430 [0.315, 0.543] -
Janowczyk 0.290 [0.259, 0.324] 0.350 [0.322, 0.376] 0.341 [0.299, 0.381] 0.135 [0.110, 0.163] 0.103 [0.088, 0.120] 0.228 [0.196, 0.259] 0.394 [0.361, 0.427] $\leq 0.01$
MIDOG 0.159 [0.136, 0.185] 0.155 [0.139, 0.172] 0.274 [0.238, 0.308] 0.342 [0.312, 0.370] 0.434 [0.400, 0.469] 0.009 [0.007, 0.011] 0.388 [0.349, 0.426] -
NuCLS 0.227 [0.211, 0.242] 0.170 [0.155, 0.186] 0.172 [0.157, 0.188] 0.347 [0.330, 0.362] 0.360 [0.343, 0.377] 0.227 [0.213, 0.243] 0.371 [0.354, 0.389] $\leq 0.01$
Overall 0.462 [0.459, 0.466] 0.472 [0.468, 0.476] 0.502 [0.499, 0.506] 0.526 [0.523, 0.529] 0.432 [0.429, 0.435] 0.242 [0.240, 0.245] 0.671 [0.668, 0.674] $\leq 0.01$
@@
Irregularity Interval MedSAM (Union Box) PathSegmentor
@@
Instance Ratio Interval MedSAM (Union Box) PathSegmentor
@@
Instance Count Interval MedSAM (Union Box) PathSegmentor
@@
Table 11: Prompt Efficiency of PathSegmentor and MedSAM. The column “#Prompts” indicates the average number of prompts used per mask, while “Dice” represents the average Dice score for that group.

### v2

@@ 2.6 PathSegmentor for Explainable Cancer Diagnosis
Explainable AI in medical image analysis is essential for enhancing the interpretability and trustworthiness of diagnostic models [36, 37, 38, 39]. In pathology, accurate cancer diagnosis relies on the identification, morphological assessment, and quantitative analysis of specific pathological objects, such as characteristic tissue patterns, cell structures, and nuclear features [40, 41]. Therefore, precise segmentation of these objects is critical for diagnostic decision-making. By extracting and analyzing diagnostically relevant biomarkers, PathSegmentor can generate interpretable, human-understandable explanations that help pathologists make more precise, evidence-based diagnostic judgments.
@@ 2.6.1 Classification-Segmentation for Feature Importance Estimation
The standard classification model for breast cancer diagnosis is trained on WSIs from the TCGA-BRCA dataset, following the conventional pipeline of patch feature extraction, slide aggregation, and final classification (Fig. 11a and Section 4.3). This model achieves high classification performance with a macro AUC of 0.936 (Supplementary Table 19). To investigate model behavior, we implement an object-based perturbation approach to estimate feature importance. Different from traditional methods that use non-semantic perturbations, such as gray squares [44] or superpixel occlusions [45], our approach aims to perform biologically meaningful perturbations. We utilize PathSegmentor to segment predefined pathological objects within an image, and then apply localized blurring perturbations exclusively to these segmented regions. Feature importance is quantified by measuring the relative increase in prediction error (Section 4.3). For example, when applying perturbations to segmented regions of breast-tissue-tumor, the model’s prediction error increases from 0.003 to 0.012 (Fig. 10a, top right). Given a set of predefined pathological objects relevant to IDC and ILC diagnosis, the analysis across all TCGA-BRCA testing samples (Fig. 10a, bottom) reveals that these pathological objects generally contribute more significantly to the IDC class than ILC. Notably, the highly discriminative IDC features discovered by PathSegmentor, including breast-nuclei-ductal epithelium, breast-cell-cancerous, breast-nuclei-neoplastic, breast-tissue-tumor, breast-nuclei-epithelial, and breast-tissue-dcis, well correspond to established pathological diagnostic criteria for the IDC class [46]. These results establish PathSegmentor as an effective explanation tool that both validates model’s learning of medically relevant patterns and provides interpretable insights by quantifying decisive features in histological terms.
@@ 2.6.2 Segmentation-Classification for Imaging Biomarker Discovery
Another approach to improve the interpretability of standard classification models for cancer diagnosis is to develop an object-aware classification model capable of generating object-aware visual explanations. In our experiments, we build a segmentation-classification framework (Fig. 10b), first leveraging PathSegmentor to predict masks of pathological objects, and then decomposing whole-image features into explicit pathological object representations, using object-based multi-instance learning to make final predictions (Fig. 11b and Section 4.3). While Class Activation Maps (CAMs) [47] from standard classification models highlight prediction-relevant regions that potentially serve as imaging biomarkers [48], their clinical utility is limited by an inability to specify the pathological nature of salient regions [49]. Our object-aware classification model bridges this semantic gap by generating object-aware CAMs, which not only localize discriminative regions but also associate them with specific pathological objects. This capability significantly enhances clinical interpretability, as pathologists require both precise pathological identification and spatial localization to discover biomarkers and inform diagnostic decisions.
After training on the TCGA-BRCA dataset, our object-aware classification model maintains diagnostic performance with a macro AUC of 0.953 for IDC and ILC classification (Supplementary Table 19). We generate object-aware CAMs by measuring the similarity between object features and classifier weights (Section 4.3). Comparing vanilla CAMs with object-aware CAMs (Fig. 10c), we observe that both methods identify similar discriminative regions indicated by warmer color intensities, while our method additionally provides semantic information about these regions. For the IDC case in the top row, our object-aware CAM accurately localizes compact tumor mass regions and, importantly, identifies imaging biomarkers with high activation values that include breast-tissue-necrosis or debris, breast-tissue-tumor, breast-nuclei-epithelial. For the ILC case in the bottom row, our method effectively captures dispersed infiltration patterns of cancerous cells and ranks the salience of pathological biomarkers, highlighting the top three as breast-cell-cancerous, breast-nuclei-ductal epithelium, breast-nuclei-neoplastic. By resolving the semantic ambiguity of traditional CAMs, PathSegmentor delivers clinically meaningful visual explanations that assist pathologists in their diagnostic processes and has the potential to uncover previously under-recognized biomarker candidates.
@@ 3 Discussion
2) PathSegmentor could benefit from integrating universal prompts. PathSegmentor’s text prompts enable efficient segmentation; however, performance may decline with semantically novel categories. Spatial prompts provide complementary localization precision. We will combine multiple prompts to enhance robustness [16, 51, 52], especially for unseen categories in complex scenarios.
3) PathSegmentor requires further real-world clinical validation. The clinical utility of PathSegmentor requires validation through multicenter trials that assess real-world segmentation accuracy and robustness. Iterative refinements of PathSegmentor will be guided by pathologist feedback to ensure seamless integration into clinical workflows.
@@ 4.1 PathSeg Dataset Construction
In this work, we create the PathSeg dataset by integrating 21 publicly available pathology image segmentation datasets, totaling 275k image-mask-label triples. The statistical details and download links of each dataset are presented in Supplementary Table 1. The diverse pathology image samples included in the PathSeg dataset cover 20 anatomical regions, 3 histological structures, and 61 object types.
[…]
@@
Data standardization. Prior to model training, we perform data preprocessing to standardize images from different datasets. The processing includes the following steps on both images and corresponding masks. 1) Magnification normalization. According to dataset metadata, images from 12 datasets in the PathSeg dataset have an initial magnification of $40\times$ (Supplementary Table 1). To standardize magnification, images from the remaining datasets are rescaled to $40\times$ using bilinear interpolation, while the corresponding masks are rescaled using nearest-neighbor interpolation. This approach ensures that the same object type maintains a relatively consistent size across different images, preventing significant variations in appearance. 2) Patching. Whole-slide images are split into a series of image patches using a sliding window. For any image dimension $D$ in height or width, if it exceeds 1500 pixels, we apply a sliding window of size 1024. The uniform overlap between adjacent patches is calculated as $\frac{(1024\times\left\lceil D/1024\right\rceil)-L}{\left\lceil D/1024\right\rceil-1}$. This ensures complete coverage while maintaining spatial regularity. The same method is applied to corresponding masks. 3) Resolution standardization. Image patches are resized to a consistent $\times1024$ resolution via bilinear interpolation, and the mask patches are resized using nearest-neighbor interpolation. This unified input dimension facilitates batch processing during training. Finally, we split the datasets for model training and testing (Supplementary Table 1). We use the official split when available. If no official split is provided, we randomly divide the original dataset into 80% for training and 20% for testing.
@@ 4.2 PathSegmentor Architecture
We develop PathSegmentor, a pathology segmentation foundation model that can handle a wide spectrum of pathological objects using textual prompts. Formally, each segmentation sample is represented as a triplet $(\textbf{x},\textbf{y},\textbf{t})$, where $\textbf{x}\in\mathbb{R}^{H\times W\times 3}$ is the input pathology image, $\textbf{y}\in\mathbb{R}^{H\times W}$ is the binary mask annotation, t is the textual prompt describing the target object. As depicted in Fig. 2a, PathSegmentor follows SEEM [16] with three key components, including an image encoder $\Phi_{\text{image}}$ that extracts visual features of the pathology image, a text encoder $\Phi_{\text{text}}$ that generates semantic embeddings from the textual prompt, and a joint feature interaction module $\Phi_{\text{joint}}$ where learnable queries fuse multi-modal features through cross-attention and self-attention mechanisms. Overall, the model predicts the object mask $\mathbf{\hat{y}}$ in the image $\mathbf{x}$ using the textual prompt $\mathbf{t}$:
@@
$\textbf{q}^{\prime}$ $=\text{MultiHead-CrossAttn}(\textbf{q},F_{\text{image}})$ (4)
@@
Second, the image-enhanced queries $\mathbf{q}^{\prime}\in\mathbb{R}^{n\times d}$ are concatenated with text features $F_{\text{text}}$ to perform multi-head self-attention. Specifically, $\mathbf{q}^{\prime}\,||\,F_{\text{text}}]\in\mathbb{R}^{(n+L)\times d}$ serve as the query, key, and value for computing the joint features $F_{\text{joint}}$, formulated as:
$F_{\text{joint}}$ $=\text{MultiHead-SelfAttn}([\mathbf{q}^{\prime}\,||\,F_{\text{text}}])$ (5)
@@
[…]
@@ 4.4 Implementation Details
Explainable cancer diagnosis. For both standard and object-based cancer diagnosis models (Fig. 11), we employ the image encoder of CONCH [28] as a frozen feature extractor, training only the slide aggregator and classifier. We adopt Attention-Based Multiple Instance Learning (ABMIL) [58] as the slide aggregator and a fully connected layer as the classifier. Following the CLAM [59] toolkit’s preprocessing pipeline, we extract patches with $\times1024$ pixels at level 0 for all slides, resizing them to $\times224$ pixels while excluding slides with insufficient tissue area. We use the Adam [60] optimizer with a learning rate of 2e-4 and a weight decay of 1e-5. The diagnosis models are implemented in PyTorch [57] and run on one NVIDIA GTX3090 GPU.
@@ 4.6 Competing Methods
All datasets in the PathSeg dataset are publicly available, with a full list of dataset links and statistical details being presented in Supplementary Table 1.
@@
The code for PathSegmentor is available at https://anonymous.4open.science/r/PathSegmentor-3166. For the competing methods, we adopt the official implementations of nnU-Net (https://github.com/MIC-DKFZ/nnUNet), DeepLabV3+ (https://github.com/VainF/DeepLabV3Plus-Pytorch), SAM-Path (https://github.com/cvlab-stonybrook/SAMPath), MedSAM (https://github.com/bowang-lab/MedSAM), SAM-Med2D (https://github.com/OpenGVLab/SAM-Med2D), and BiomedParse (https://github.com/microsoft/BiomedParse).
Ethics Declarations
This project has been reviewed and approved by the Human and Artefacts Research Ethics Committee (HAREC) of The Hong Kong University of Science and Technology. The protocol number is HREP-2025-0024.
@@ Acknowledgment
Appendix A Supplementary Material
Table 1: An overview of the 21 selected datasets in our PathSeg dataset for pathology image segmentation. The table lists dataset name, magnification (MA), anatomical regions (AR), number of labels, images, masks (train/test masks).
@@ Appendix A Supplementary Material
Internal Datasets (16)
BCSS [64] 40x Breast 18 2,914 9,192 (7,322/1,870)
CoCaHis [65] 40x Liver 1 82 82 (58/24)
CoNIC2022 [66] 20x Colon 6 4,841 19,332 (15,387/3,945)
CRAG [67] 20x Colon 1 1,750 1,750 (1,429/321)
DigestPath2019 [68] 40x Colon 1 13,391 13,391 (10,725/2,666)
GlaS [69] 20x Colon 1 165 165 (85/80)
Janowczyk [70] 40x Breast 1 500 500 (400/100)
MIDOG [71] 40x Unspecified 2 1,482 1,802 (1,448/354)
MoNuSAC [72] 40x Unspecified 4 359 642 (454/188)
NuCLS [73] 40x Breast 12 1,612 4,443 (3,538/905)
PanNuke [74] 40x Unspecified 85 7,558 16,223 (13,017/3,206)
RINGS [75] 100x Prostate 2 38,227 43,739 (29,451/14,288)
SegPath [76] 40x Unspecified 8 133,645 133,645 (122,536/11,109)
SICAPv2 [77] 20x Prostate 1 23,924 23,924 (19,140/4,784)
TNBC [78] 40x Breast 11 66 167 (134/33)
WSSS4LUAD [79] 10x Lung 3 2,863 5,735 (4,590/1,145)
External Datasets (5)
CPM17 [80] 20x, 40x Unspecified 1 15 15 (0/15)
CPM15 [80] 20x, 40x Unspecified 1 64 64 (0/64)
Kumar [81] 40x Unspecified 1 238 238 (0/238)
Lizard [82] 20x Unspecified 1 30 30 (0/30)
CoNSeP [8] 40x Colon 6 40 121 (0/121)
@@
DeepLabV3+ SAM-Path nnU-Net
MedSAM
(Union Box)
SAM-Med2D
(Union Box)
BiomedParse PathSegmentor
GlaS
0.735
[0.702, 0.765]
0.833
[0.816, 0.850]
0.905
[0.885, 0.922]
0.843
[0.814, 0.868]
0.650
[0.621, 0.675]
0.862
[0.842, 0.880]
0.937
[0.924, 0.947]
$\leq 0.01$
CRAG
0.752
[0.733, 0.772]
0.814
[0.798, 0.828]
0.895
[0.878, 0.911]
0.765
[0.746, 0.783]
0.583
[0.561, 0.604]
0.559
[0.531, 0.588]
0.933
[0.925, 0.940]
$\leq 0.01$
WSSS4LUAD
0.677
[0.661, 0.692]
0.755
[0.742, 0.767]
0.838
[0.825, 0.850]
0.405
[0.391, 0.419]
0.379
[0.365, 0.392]
0.036
[0.032, 0.039]
0.833
[0.822, 0.843]
CoCaHis
0.663
[0.582, 0.736]
0.757
[0.704, 0.806]
0.822
[0.755, 0.881]
0.611
[0.518, 0.692]
0.509
[0.388, 0.609]
0.679
[0.615, 0.734]
0.827
[0.791, 0.861]
$0.06$
SICAPv2
0.608
[0.597, 0.618]
0.682
[0.674, 0.691]
0.622
[0.612, 0.632]
0.806
[0.801, 0.811]
0.700
[0.694, 0.706]
0.241
[0.235, 0.246]
0.821
[0.813, 0.828]
$\leq 0.01$
RINGS
0.602
[0.596, 0.608]
0.733
[0.728, 0.739]
0.661
[0.655, 0.667]
0.775
[0.771, 0.779]
0.585
[0.582, 0.589]
0.246
[0.243, 0.248]
0.800
[0.794, 0.805]
$\leq 0.01$
DigestPath2019
0.643
[0.631, 0.656]
0.539
[0.525, 0.554]
0.661
[0.650, 0.674]
0.764
[0.757, 0.772]
0.648
[0.640, 0.657]
0.508
[0.498, 0.518]
0.784
[0.775, 0.794]
$\leq 0.01$
CoNIC2022
0.461
[0.452, 0.470]
0.360
[0.351, 0.368]
0.650
[0.642, 0.658]
0.202
[0.194, 0.211]
0.169
[0.163, 0.176]
0.287
[0.279, 0.296]
0.681
[0.674, 0.689]
$\leq 0.01$
MoNuSAC
0.534
[0.493, 0.573]
0.493
[0.456, 0.533]
0.644
[0.605, 0.682]
0.359
[0.323, 0.400]
0.347
[0.310, 0.387]
0.288
[0.240, 0.336]
0.657
[0.615, 0.698]
$\leq 0.01$
PanNuke
0.541
[0.530, 0.551]
0.336
[0.324, 0.347]
0.564
[0.553, 0.574]
0.364
[0.355, 0.374]
0.307
[0.298, 0.317]
0.621
[0.612, 0.631]
0.631
[0.621, 0.641]
$\leq 0.01$
BCSS
0.393
[0.377, 0.408]
0.461
[0.445, 0.478]
0.506
[0.488, 0.523]
0.568
[0.555, 0.581]
0.508
[0.496, 0.521]
0.126
[0.117, 0.135]
0.620
[0.605, 0.636]
$\leq 0.01$
SegPath
0.162
[0.157, 0.167]
0.104
[0.100, 0.108]
0.122
[0.117, 0.126]
0.215
[0.210, 0.219]
0.194
[0.190, 0.199]
0.083
[0.080, 0.085]
0.442
[0.436, 0.447]
$\leq 0.01$
TNBC
0.209
[0.130, 0.296]
0.251
[0.137, 0.379]
0.473
[0.350, 0.593]
0.260
[0.184, 0.348]
0.229
[0.160, 0.306]
0.361
[0.243, 0.480]
0.430
[0.315, 0.543]
Janowczyk
0.290
[0.259, 0.324]
0.350
[0.322, 0.376]
0.341
[0.299, 0.381]
0.135
[0.110, 0.163]
0.103
[0.088, 0.120]
0.228
[0.196, 0.259]
0.394
[0.361, 0.427]
$\leq 0.01$
MIDOG
0.159
[0.136, 0.185]
0.155
[0.139, 0.172]
0.274
[0.238, 0.308]
0.342
[0.312, 0.370]
0.434
[0.400, 0.469]
0.009
[0.007, 0.011]
0.388
[0.349, 0.426]
NuCLS
0.227
[0.211, 0.242]
0.170
[0.155, 0.186]
0.172
[0.157, 0.188]
0.347
[0.330, 0.362]
0.360
[0.343, 0.377]
0.227
[0.213, 0.243]
0.371
[0.354, 0.389]
$\leq 0.01$
Overall
0.462
[0.459, 0.466]
0.472
[0.468, 0.476]
0.502
[0.499, 0.506]
0.526
[0.523, 0.529]
0.432
[0.429, 0.435]
0.242
[0.240, 0.245]
0.671
[0.668, 0.674]
$\leq 0.01$
@@
Irregularity
Interval
MedSAM (Union Box) PathSegmentor
@@
Instance Ratio
Interval
MedSAM (Union Box) PathSegmentor
@@
Instance Count
Interval
MedSAM (Union Box) PathSegmentor
@@
Table 11: Prompt Efficiency of PathSegmentor and MedSAM. The column of #Prompts indicates the average number of prompts used per mask, while Dice represents the average Dice score for that group.